OpenStudy (decol):

Given a geometric sequence with a1 = 6 and r = 2/3, write an explicit formula for an, the nth term of the sequence.

OpenStudy (solomonzelman):

The recursive formula for the sequence is in a form of, \(\color{#000000 }{ \displaystyle a_n=a_1 \times r^{n-1} }\)

OpenStudy (solomonzelman):

You are given the first term \(\color{#000000 }{ \displaystyle a_1 }\), and the common ratio \(\color{#000000 }{ \displaystyle r }\). Go for it!

OpenStudy (decol):

Thanks

OpenStudy (solomonzelman):

YW